OpenStudy (anonymous):

Let A=(0,0) and B=(12,5). Find points on the y-axis that are twice as far from B as from A. Solve algebraically.

OpenStudy (campbell_st):

ok, this problem and needs the distance formula call the point on the y axis P(0, y) then you have 2AP = BP \[2\sqrt{(0 - 0)^2 + (y - 0)^2} = \sqrt{(0 - 12)^2 + ( y - 5)^2}\] square both sides \[4y^2 = 144 + (y - 5)^2\] you need to now solve for y

OpenStudy (campbell_st):

the solution will need the general quadratic formula.

OpenStudy (anonymous):

Thank you!